Shift report 20011123 08:00-16:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Fri Nov 23 2001 - 00:10:24 EST

  • Next message: Trine S. Tveter: "Re: width of dnd/deta"

    ------------------------------------------------------------
    Date: Fri Nov 23 00:10:23 2001, Shift: 08:00-16:00
    Supervisor: James Norris, with: 
    Log book#11 pages 25 - 26
    
    Shift Summary:
    ------------------------------------------------------------
    1. At 16:00 previous beam stored & cogged 
          (at ~15:00 22-Nov-01)
       At 17:00 switched D1-D4 to polarity B for the 
                FS=8degree D1-D4 1/8B Run
       At 17:15 MCR announced 15min access at end of this run
       At 19:55 voted to dump
       At 19:55 MCR announced beam dump in 30-45min
       At 20:21 MCR called to warn of beam dump 15min
       At 20:50 beam was dumped
       At 21:40 MCR estimates beam in 30min
       At 22:15 MCR preparing to fill RHIC
       At 22:47 beam stored and cogged, ZDC rates:
          BRAHMS=550, STAR=1351, PHENIX=2598, PHOBOS=1578
    
                                 trigger 6
    2. Run 5947 - 15:50-16:57      102.3K
       Run 5947 completed the FS=8degree D1-D4 1/8A Goal
       Run 5948 - 17:21-20:30      105.8K
       Run 5952 - 22:55-23:23       55.6K
       Run 5952 completed the FS=8degree D1-D4 1/8B Goal
       Run 5953 - 23:29-           (70.1K)
    
    3. Nada!
    
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b30 : Fri Nov 23 2001 - 00:10:55 EST